OpenStudy (anonymous):

State the slope and the y-intercept for the graph of the equation 5x+3y=0 please help .. im so confusedd..

OpenStudy (anonymous):

first put what you have in slope intercept form

OpenStudy (anonymous):

i dont even know what that means...

OpenStudy (anonymous):

y= mx+b were m is the slope and b is the y-intercept

OpenStudy (dan815):

put it in terms of Y and then we will talk

OpenStudy (anonymous):

@katelm123 so m=3y and = 0 ?

OpenStudy (anonymous):

first solve for y

OpenStudy (anonymous):

i dont know how ??

OpenStudy (anonymous):

Taliea, first you want to separate the x and y. Can you do that?

OpenStudy (anonymous):

if i knew what the x and y were i would ?

OpenStudy (anonymous):

5x is the "x" and 3y is the "y"

OpenStudy (anonymous):

5x+3y=0 -5x -5x 3y=-5x dive by 3 on both sides y= -5/3x

OpenStudy (anonymous):

so the slope is -5/3 and the y intercept is 0

OpenStudy (anonymous):

@katelm123 I don't think she understands that yet, we need to take it slower.

OpenStudy (anonymous):

ok lets see...

OpenStudy (anonymous):

oh @katelm123 i kinda understad from you working it out :)

OpenStudy (dwelsher):

1.put equation in slope intercept form first 2 y=mx+b 3. m=slope , b equals y intercepy

OpenStudy (dan815):

^

OpenStudy (anonymous):

Oh okay, I was getting worried about how to teach you it xD So do you understand everything now?

OpenStudy (anonymous):

yes i do thanks :D

OpenStudy (dan815):

i dont trust you do this! 5x-2y=4

OpenStudy (anonymous):

Haha, I agree, for your sake show us you can do that.

OpenStudy (dwelsher):

3y=-5x+o y=-5/3x+0/3 ,y=-5/3X+0 WHERE ARE YOU M AND CONSTANT?

OpenStudy (dan815):

stahhhp dwel i have a vague suspicion if she can isolate for y or not

OpenStudy (anonymous):

dan your kind of annoying but the answer is 5/2?

OpenStudy (dan815):

no -_-

OpenStudy (anonymous):

Aww but Dan is trying to help you learn it. And it is not 5/2 =P

OpenStudy (anonymous):

dan is doing the correct thing

OpenStudy (anonymous):

i tried -.- math is hard ok .

OpenStudy (anonymous):

Show us your work, I'm sure we can find the problem.

OpenStudy (dan815):

draw like this